OpenStudy (lgbasallote):

Note: This is NOT a question. This is a tutorial. how to factor quadratic expressions? For example you are given \(x^2 + 3x - 4\). See comment below to see how

OpenStudy (lgbasallote):

The first step would be to break down the middle term into addends such that the first addend is a factor of the first term (in this case we are given \(x^2\) as the first term), and the secod addend must be a factor of the last term (in this case we are given \(-4\) as the last term). Now, let us think of addends of +3x. Since our first term is positive, and our last term is negative, then our addends must have one positive, and one negative. The first trial would be +x and -4x. So let's try it. \(x^2 + 3x - 4 = x^2 - x + 4x - 4\) from here, we can clearly see that \((x^2 - x)\) is factorable, and \((4x - 4)\) is too. So let's factor them \(x(x - 1)\) is the factored form of \((x^2 -x)\) \(4(x-1)\) is the factored form of \((4x - 4)\) So combining these data, we have... \(x(x - 1) + 4(x - 1)\) Notice how both terms have (x - 1). Therefore, we can say that (x-1) can be factored. If we do that, we'll have.. \((x-1)(x+4)\) This is now the factored form of \(x^2 + 3x - 4\)

Parth (parthkohli):

Why'd you ask haha?

OpenStudy (lgbasallote):

because im bored...im doing tutorials :D

Parth (parthkohli):

Good idea :)

OpenStudy (lgbasallote):

lol my first paragraph had something wrong. "...The first trial would be +x and -4x" it should be -x and +4x. Please note this correction

OpenStudy (mimi_x3):

There is an easier way.. \[a+b=-3\] \[ab=-4\] \[a = -4; b=1\] so.. \[(x-4)(x+1)\]

OpenStudy (anonymous):

(x+a)(x+b) = x^2 + x(a+b) + ab so you want 2 numbers...etc Need to explain what if coefficient of x^2 is not 1 (Assuming "easy" factors)

OpenStudy (lgbasallote):

i am not aware of this @Mimi_x3 o.O is that a real theorem? i mean is it ripped off from a real one (like binomial theorem or something)

OpenStudy (mimi_x3):

lol, i can't remember what it is called..I tend to do it in my head. How can it be the Binomial Theorem? lol

Parth (parthkohli):

\[\LARGE x^{2} + 3x - 4\] \[\LARGE x^{2} -x + 4x - 4\] \[\LARGE x(x - 1) + 4(x - 1) \] \[\LARGE (x + 4)(x - 1) \] \[\LARGE \text {EASY, BABY ;)} \]

OpenStudy (lgbasallote):

lol i do not know Mimi :p but at least i learned a new one today :) this discussion was a good idea after all good work @ParthKohli :D

OpenStudy (mimi_x3):

Want a Binomial Theorem example?

OpenStudy (lgbasallote):

NO!

OpenStudy (mimi_x3):

Its fun!!

OpenStudy (lgbasallote):

@estudier 's algebraic equation is equivalent to the one you said mimi?

OpenStudy (lgbasallote):

yeah right fun you sadistic bunny :P

OpenStudy (mimi_x3):

c'mon you will love it :P

Parth (parthkohli):

@lgbasallote Please edit the question. It's how to factor quadratic expressions, not equations. :)

OpenStudy (lgbasallote):

thankk you for the correction :) and @Mimi_x3 i'll take permutation combination and probability but binomial theorem..NO WAY!

OpenStudy (callisto):

 That's how I did it